### v3

@@ Abstract
Do social networks and peer influence shape major life decisions in highly polarized settings? We explore this question by examining how peers influenced the allegiances of West Point cadets during the American Civil War. Leveraging quasi-random variations in the proportion of cadets from Free States, we analyze how these differences affected cadets’ decisions about which army to join. We have three main findings. First, there was a strong and significant peer effect: a higher proportion of classmates from Free States significantly increased the likelihood that cadets from Slave States joined the Union Army. Second, the peer effect varies with geography, most notably with the slave population share in cadets’ home states or counties, and with cadets’ own slave ownership in 1860. Third, shared experiences—such as having served together in the Mexican-American War, continuous military service, and belonging to the same cohort—amplified peer effects, suggesting that sustained interaction is important.
@@ 1 Introduction
Do social networks and peer influence shape individuals’ choices when societies are deeply polarized? This question is central not only to understanding moments of historical rupture but also to how political allegiances are formed in contemporary democracies. The literature on polarization has highlighted partisan sorting, media biases, divergent economic situations, and cultural divisions as key drivers of political alignment (e.g., Mason, 2018; Prior, 2013; Autor et al., 2020; Hochschild, 2016). Less is known about the role of peers and social networks in shaping political loyalties in polarized settings. While research has shown that peers matter for outcomes such as education, consumption, health, and careers, [^1] We cannot include a comprehensive review of the literature here, but such overviews are available in Benhabib et al. (2010), Brechwald and Prinstein (2011), Aral and Nicolaides (2017), McGloin and Thomas (2019), Giletta et al. (2021), and Boucher et al. (2024). there is little causal evidence as to whether peers influence high-stakes political and/or life decisions, especially when these decisions come with severe personal risks and moral conflicts.
@@
The American Civil War provides a unique opportunity to examine this question. West Point cadets decisions as to which side to fight for was often the most consequential choice of their lives, pitting personal backgrounds, political-economic interests, and regional loyalties against professional ties and peer influences. Although West Pointers played pivotal roles in the war, systematic empirical analysis of the factors influencing their allegiances remains limited. [^2] There exists a rich historical literature on West Pointers in the Civil War (e.g., Patterson, 2002). A recent political science study, White (2024), highlights the importance of economic interests. See also additional references below. This setting offers two key advantages. First, the quasi-random composition of cadet cohorts generates exogenous variation in exposure to peers from different regions. This allows us to separate peer effects from confounding factors (e.g., individuals typically self-select into like-minded groups) that typically make it difficult to disentangle network influence from shared beliefs or communal forces (see Manski, 1993; Angrist, 2014). Quasi-randomness arose from institutional features: although admissions broadly tracked the geographic distribution of congressional districts, annual recruitment varied substantially due to decentralized nomination procedures and qualification exams (e.g., Park, 1840). Second, West Point was the nation’s premier military academy, where cadets from Northern (Free) and Southern (Slave) lived and trained together, and forged close bonds before the war. This environment magnified the potential for peer influence. Historians have long noted that “group cohesion” and camaraderie motivated enlistment and fighting, but the evidence has been largely anecdotal, drawn from letters and diaries (e.g., McPherson, 1997, 2003; Siebold, 2007). Whether Northern peers swayed Southern cadets toward the Union, or whether local political and economic pressures outweighed such influence, has been an open question.
@@
The framework also implies that peer effects vary with political-economic background, which shapes the tradeoff inherent in the allegiance decision. Using slave-holding data from the 1860 Census and genealogical records from Ancestry.com and Familysearch.com, we find consistent evidence for this prediction. Among cadets from Slave States, stronger ties to slavery, proxied by a very high slave population shares in one’s home state (or county) or by personal slave ownership, essentially eradicate peer influence. The peer effect is very large in regions and cadets less tied to slavery. Moreover, slave population share emerges as a more powerful proxy for political-economic background than alternative measures, including county-level presence of pro- or anti-slavery religious groups and county-level voting shares for pro- and anti-slavery candidates in the 1860 presidential election.
Our third set of analyses highlights the central role of interaction in shaping peer influence. We find that peer effects are substantially stronger within cohorts than across cohorts, indicating the importance of direct and contemporaneous interaction. Peer influence is also more pronounced among cadets who maintained continuous military careers between their graduation and the Civil War, compared with those who had returned to civilian life, suggesting that sustained institutional embeddedness amplified interpersonal influence. We also track variation in cadets’ likelihood of having served together in the Mexican–American War (1846–1848). Peer effects are stronger when cadets from Slave States had more Free-State peers with whom they shared combat experience, and weaker when their co-combatants were also from Slave States. Together, these findings provide additional evidence that peer influence operates through repeated/sustained interactions.
@@
Our study contributes knowledge about peer effects (see Benhabib et al. (2010); Brechwald and Prinstein (2011); McGloin and Thomas (2019); Giletta et al. (2021) for some overviews), by demonstrating how peer influence can shape major life decisions, particularly during moments of extreme polarization. More broadly, it adds to the long-standing literature emphasizing the importance of “personal influence” in economic, political, and social behaviors (e.g., Katz and Lazarsfeld, 1955). Other research specifically involving soldiers concerns the spread of political ideas (Jha and Wilkinson, 2023; Grosjean et al., 2025), which our study complements in studying allegiance (and how it depends on economic circumstances).
@@
The experiences of West Point cadets during the American Civil War mirror other historical contexts where military academies produced leaders for both sides of major conflicts in divided nations. For instance, the Whampoa Military Academy in the 1920s produced commanders who fought on opposing sides during the Chinese Civil War (e.g., Bai et al., 2023a), and the Heroic Military Academy in Mexico trained leaders who both supported and opposed the regime during the Mexican Revolution. Such decisions are not unique to military cadets, as similar choices are faced by ordinary citizens in many divided societies. However, analyzing peer influence in these broader contexts is often more challenging. The distinctive setting of the American Civil War, where allegiances were closely tied to geography, provides a unique opportunity to examine peer influence with a level of identification that is difficult to achieve in other cases.
@@
Our findings also add to the vast body of historical and social sciences literature that examines loyalty decisions during American Civil War. McPherson (1997) draws on soldiers’ letters and diaries to investigate soldiers’ motivations. Analyzing Union Army company data, Costa and Kahn (2003, 2008) emphasize the importance of loyalty to fellow soldiers, which sometimes surpassed commitment to the cause when considering decisions to fight, shirk duty, or desert. Our findings suggest that both cause (the allegiance to one’s home state) and comrades (peer influence) matter and that they also interact with each other. Closely related, Hall et al. (2019) document that slave ownership drove enlistment in the Confederacy, whereas White (2024) explores the allegiance of West Pointers, emphasizing economic interest by showing that one’s past employment related to cash crops (i.e., sugar, indigo, rice, tobacco, and cotton) is linked to the likelihood of West Pointers joining the Confederacy. [^3] Using the same variable on cash-crop employment, we find a similar pattern. But with only a small share of candidates having such a history, we do not find a significant interaction effect between this variable and peer influence. Our observation regarding the influence of the slave economy corroborates these findings, though our primary focus remains on peer influence—a critical factor in loyalty decisions during pivotal historical moments.
@@
Beyond this context, a large literature has examined the determinants of conflict participation, highlighting the importance of interests (as reviewed in Blattman and Miguel (2010)) and social forces (reviewed in Cantoni et al. (2024)). [^4] There is also a long-standing literature in sociology recognizing the importance of personal networks in social and political movements (e.g., McAdam, 1986, Diani and McAdam, 2003). We contribute to the previous literature in terms of the variation in peer composition that identifies peer influence, the differing ways that it works depending on background of a cadet’s state, the decisions made by leaders, and the choice between competing ideologies.
@@
Finally, our study complements an expanding literature on leadership (e.g., Jones and Olken, 2005; Acemoglu and Jackson, 2015; Dippel and Heblich, 2021; Cagé et al., 2023; Bai et al., 2023b; Ferrara et al., 2025). Understanding how leaders develop their ideas and loyalties is central to grasping important political and economic outcomes. Our findings highlight the significance of peer influence in shaping future leaders’ decisions during crises, demonstrating how peer influence can drive individual choices that may shape the course of history.
@@ 2.1 The Cadets
We focus on cadets who graduated from West Point Military Academy between 1820 and 1860.555We begin in 1820 because we observe that some cadets from the earlier cohorts participated in the Civil War. We also report additional results that restrict the sample to younger cohorts only. We manually collected information from the comprehensive Biographical Register of the Officers and Graduates of the United States Military Academy (Cullum, 1891). The records provide details such as the year of graduation, birth year, home state and appointment state, [^6] Home state and appointment state usually overlap. We use home state in our main analysis and present additional robustness checks using appointment state to define peers in the Supplemental Appendix. and academic rank. To ensure comparability of academic performance, we classify cadets into percentiles based on their graduation ranking. The top 10% were assigned a score of 100, the bottom 10% a score of 10, and so on. Additionally, the records include information on wars and battles one participated in and their military ranks, which we use to determine if one joined the Union or Confederate forces during the Civil War. We present examples of these records and our coding procedure in Supplemental Appendix A.1.
@@
We collected additional data on cadet backgrounds including (i) a cadet’s home county, which can be linked to county-level characteristics; (ii) Cadet slave ownership (whether the cadet personally owned slaves in 1860); (iii) whether a cadet’s father and mother originated from Free States; and (iv) whether a cadet’s wife originated from a Free State (with unmarried cadets coded as not having such a spouse). These data are obtained by Ancestry.com and FamilySearch.com, including linked data from the 1860 Census. We describe our data-linking procedure in Supplemental Appendix A.1.
@@ 3.1 Descriptive Evidence
Note. This figure plots the relationship between the share of the slave population in a state and the likelihood that cadets from that state would join the Union. Panel A concentrates on the states that participated in the war, while Panel B uses all cadets, including those who did not participate in the war. The circle’s size represents the number of West Point cadets hailing from each state.
@@ 3.2 Statistical Analysis
$X_{i,s}$ includes a range of personal characteristics. Among them, our main analysis focuses on variables available for all cadets recorded in (Cullum, 1891): age in 1860, academic rank at West Point, and the slave population share in the cadet’s home state. When home-state fixed effects are included, state-level characteristics are naturally absorbed. In addition, we consider alternative measures of political-economic background, including whether a cadet owned slaves in 1860, and indicators for whether the cadet’s father, mother, or wife originated from Free States. These variables are constructed for the subset of cadets for whose information is available from Ancestry.com and FamilySearch.com.
@@
Our specification examines peer effects based on exogenous peer characteristics and not endogenous peer choices (Manski (1993)). Given that cadets’ decisions were made some years after graduation, it is less likely that there was direct coordination on decisions rather than lasting impacts of interactions from their formative years. Nonetheless, it is possible that there is correlation in their decisions, and thus for robustness, we also re-estimate the model using standard errors clustered at the cohort level as well as bootstrapped standard errors clustered at the cohort level. We also directly examine residual dependence across peer groups by estimating intra-class correlation coefficients (ICC) for the baseline residuals within cohorts. The results show negligible within-cohort residual correlation, providing no evidence of jointly determined wartime allegiance decisions.
@@ 3.3 The Impact of Peers on Cadets’ Allegiance
Note. This table presents the impact of the fraction of peers from Free States in a cadet’s cohort on that cadet’s decision to join the Union. Columns (1)-(4) focus on cadets from Slave States (i.e., with a slave population higher than 1%) and Columns (5)-(8) on those from Free States (i.e., with a slave population lower than 1%). The variable Free-State father and Free-State mother equal 1 if a parent is from a Free State, and 0 otherwise. A subset of observations contains missing information on parental origin. Free-State wife equals 1 if the wife is from a Free State and 0 otherwise. If the wife’s information is missing, we assign a value of 0. The standard errors presented in the parentheses are obtained through bootstrapping with 400 resampling iterations. ***$p<0.01$, **$p<0.05$, *$p<0.1$.
We include cadets who graduated between 1820 and 1860 in our baseline analysis because even some cadets from the early cohorts (from 1820) participated in the Civil War, often serving in officer or administrative roles. As a robustness check, we re-estimate our baseline specifications after excluding those aged 60 (or 50) and above; the estimated peer effects remain quantitatively similar and statistically robust. The results are reported in Appendix Table S7.
@@
For Free-State cadets $S_{i}$ and $V$ is likely to be high, so this probability is close to 1. For cadets heavily exposed to slavery (either from a region heavily dependent on slavery, or personally owning slaves), $S_{i}$ is large and this probability is close to 0. It is for intermediate values of $S_{i}-V$ that this varies substantially as a function of $\beta F_{i}$.
@@
We further divide the Slave States into groups by the relative slave populations. It turns out that the peer influence operates entirely on cadets from the middle range of Slave States that have slave populations below one third of the population. In Panel B of Table 2, we break states into three groups: those without slaves, those with slaves up to one third of the population, and those with slave populations above one third. We now see that the peer effect comes entirely from this middle group of states, and in fact the estimated coefficient and significance are larger than that from the previous table. These findings suggest that in High-Slavery States, economic imperatives and entrenched social norms overrode peer influence, muting its effect. This is consistent with a very high $S_{i}$ in equation (2).

### v4

@@ Abstract
Do social networks and peer influence shape major life decisions in highly polarized settings? We explore this question by examining how peers influenced the allegiances of West Point cadets during the American Civil War. Leveraging quasi-random variations in the proportion of cadets from Free States, we analyze how cadets’ decisions about which army to join depended on the composition of their peers. We have three main findings. First, there was a strong and significant peer effect: a higher proportion of classmates from Free States significantly increased the likelihood that cadets from Slave States joined the Union Army. Second, the peer effect varies with geography, most notably with the slave population share in cadets’ home states or counties, and with cadets’ own slave ownership in 1860. Third, peer effects were amplified by shared experiences such as having served together in the Mexican-American War, continuous military service, and belonging to the same cohort, suggesting that sustained interaction is important.
@@ 1 Introduction
Do social networks and peer influence shape individuals’ choices when societies are polarized? This question is central not only to understanding moments of historical rupture but also to how political allegiances are formed in contemporary democracies. The literature on polarization has highlighted partisan sorting, media biases, divergent economic situations, and cultural divisions as key drivers of political alignment (e.g., Mason, 2018; Prior, 2013; Autor et al., 2020; Hochschild, 2016). Less is known about the role of peers and social networks in shaping political loyalties in polarized settings. While research has shown that peers matter for outcomes such as education, consumption, health, and careers, [^1] We cannot include a comprehensive review of the literature here, but such overviews are available in Benhabib et al. (2010), Brechwald and Prinstein (2011), Aral and Nicolaides (2017), McGloin and Thomas (2019), Giletta et al. (2021), and Boucher et al. (2024). there is little causal evidence as to whether peers influence high-stakes political and/or life decisions, especially when these decisions come with severe personal risks and moral conflicts.
@@
The American Civil War provides an opportunity to examine this question. West Point cadets decisions as to which side to fight for pitted personal backgrounds, political-economic interests, and regional loyalties against professional ties and peer influences. Although West Pointers played pivotal roles in the war, systematic empirical analysis of the factors influencing their allegiances remains limited. [^2] There exists a rich historical literature on West Pointers in the Civil War (e.g., Patterson, 2002). A recent political science study, White (2024), highlights the importance of economic interests. See also additional references below. This setting offers two key advantages. First, the quasi-random composition of cadet cohorts generates exogenous variation in exposure to peers from different regions. This allows us to separate peer effects from confounding factors (e.g., individuals typically self-select into like-minded groups) that typically make it difficult to disentangle network influence from shared beliefs or communal forces (see Angrist, 2014). Quasi-randomness arose from institutional features: although admissions broadly tracked the geographic distribution of congressional districts, annual recruitment varied substantially due to decentralized nomination procedures and qualification exams (e.g., Park, 1840). Second, West Point was the nation’s premier military academy, where cadets from Northern (Free) and Southern (Slave) lived and trained together, and forged close bonds before the war. This environment allowed for peer influence. Historians have long noted that “group cohesion” and camaraderie motivated enlistment and fighting, but the evidence has been largely anecdotal, drawn from letters and diaries (e.g., McPherson, 1997, 2003; Siebold, 2007). Whether peers swayed cadets’ choices of which army to join, or whether local political and economic pressures outweighed such influence, has been an open question.
@@
We also find that peer effects vary with political-economic background, which shapes the tradeoff inherent in the allegiance decision. Using slave-holding data from the 1860 Census and genealogical records from Ancestry.com and Familysearch.com, we find that among cadets from Slave States, stronger ties to slavery, proxied by a very high slave population shares in one’s home state (or county) or by personal slave ownership, essentially eradicate peer influence. The peer effect is very large in regions and for cadets less tied to slavery. Moreover, slave population share emerges as a more significant proxy for political-economic background with regards to cadet’s decisions than alternative measures, including county-level presence of pro- or anti-slavery religious groups and county-level voting shares for pro- and anti-slavery candidates in the 1860 presidential election.
Our third set of analyses highlights the role of interaction in shaping peer influence. We find that peer effects are stronger within cohorts than across cohorts, indicating the importance of direct and contemporaneous interaction. Peer influence is also more pronounced among cadets who maintained continuous military careers between their graduation and the Civil War, compared with those who had returned to civilian life, suggesting that sustained embeddedness amplified interpersonal influence. We also track variation in cadets’ likelihood of having served together in the Mexican–American War (1846–1848). Peer effects are stronger when cadets from Slave States had more Free-State peers with whom they shared combat experience, and weaker when their co-combatants were also from Slave States. Together, these findings provide suggestive evidence that peer influence operates through repeated/sustained interactions.
@@
Our study contributes knowledge about peer effects (see Benhabib et al. (2010); Brechwald and Prinstein (2011); McGloin and Thomas (2019); Giletta et al. (2021) for some overviews), by demonstrating how peer influence can shape major life decisions, particularly during moments of polarization. More broadly, it adds to the long-standing literature emphasizing the importance of “personal influence” in economic, political, and social behaviors (e.g., Katz and Lazarsfeld, 1955). Other research specifically involving soldiers concerns the spread of political ideas (Jha and Wilkinson, 2023; Grosjean et al., 2025), which our study complements in studying allegiance (and how it depends on economic circumstances).
@@
The experiences of West Point cadets during the American Civil War mirror other historical contexts where military academies produced leaders for both sides of major conflicts in divided nations. For instance, the Whampoa Military Academy in the 1920s produced commanders who fought on opposing sides during the Chinese Civil War (e.g., Bai et al., 2023a), and the Heroic Military Academy in Mexico trained leaders who both supported and opposed the regime during the Mexican Revolution. Such decisions are not unique to military cadets, as similar choices are faced by ordinary citizens in many divided societies. However, analyzing peer influence in these broader contexts is often more challenging. The distinctive setting of the American Civil War, where allegiances were closely tied to geography, provides an opportunity to examine peer influence with a level of identification that is difficult to achieve in other cases.
@@
Our findings also add to the body of historical and social sciences literature that examines loyalty decisions during American Civil War. McPherson (1997) draws on soldiers’ letters and diaries to investigate soldiers’ motivations. Analyzing Union Army company data, Costa and Kahn (2003, 2008) emphasize the importance of loyalty to fellow soldiers, which sometimes surpassed commitment to the cause when considering decisions to fight, shirk duty, or desert. Our findings suggest that both cause (the allegiance to one’s home state) and comrades (peer influence) matter and that they also interact with each other. Closely related, Hall et al. (2019) document that slave ownership drove enlistment in the Confederacy, whereas White (2024) explores the allegiance of West Pointers, emphasizing economic interest by showing that one’s past employment related to cash crops (i.e., sugar, indigo, rice, tobacco, and cotton) is linked to the likelihood of West Pointers joining the Confederacy. [^3] Using the same variable on cash-crop employment, we find a similar pattern. But with only a small share of candidates having such a history, we do not find a significant interaction effect between this variable and peer influence. Our observation regarding the influence of the slave economy corroborates these findings, though our primary focus remains on peer influence—a critical factor in loyalty decisions during pivotal historical moments.
@@
Beyond this context, a literature has examined the determinants of conflict participation, highlighting the importance of interests (as reviewed in Blattman and Miguel (2010)) and social forces (reviewed in Cantoni et al. (2024)). [^4] There is also a long-standing literature in sociology recognizing the importance of personal networks in social and political movements (e.g., McAdam, 1986, Diani and McAdam, 2003). We contribute to the previous literature in terms of the variation in peer composition that identifies peer influence, the differing ways that it works depending on background of a cadet’s state, the decisions made by leaders, and the choice between competing ideologies.
@@
Finally, our study complements a literature on leadership (e.g., Jones and Olken, 2005; Acemoglu and Jackson, 2015; Dippel and Heblich, 2021; Cagé et al., 2023; Bai et al., 2023b; Ferrara et al., 2025). Understanding how leaders develop their ideas and loyalties can have political and economic implications. Our findings highlight the significance of peer influence in shaping future leaders’ decisions during crises, demonstrating how peer influence drives individual choices that can have historical implications.
@@ 2.1 The Cadets
We focus on cadets who graduated from West Point Military Academy between 1820 and 1860.555We begin in 1820 because we observe that some cadets back to this date participated in the Civil War. We also report additional results that restrict the sample to later cohorts only. We manually collected information from the comprehensive Biographical Register of the Officers and Graduates of the United States Military Academy (Cullum, 1891). The records provide details such as the year of graduation, birth year, home state and appointment state, [^6] Home state and appointment state usually overlap. We use home state in our main analysis and present additional robustness checks using appointment state to define peers in the Supplemental Appendix. and academic rank. To ensure comparability of academic performance, we classify cadets into percentiles based on their graduation ranking. The top 10% were assigned a score of 100, the bottom 10% a score of 10, and so on. Additionally, the records include information on wars and battles one participated in and their military ranks, which we use to determine if one joined the Union or Confederate forces during the Civil War. We present examples of these records and our coding procedure in Supplemental Appendix A.1.
@@
We collected additional data on cadet backgrounds including (i) a cadet’s home county, which can be linked to county-level characteristics; (ii) cadet slave ownership (whether the cadet personally owned slaves in 1860); (iii) whether a cadet’s father and mother originated from Free States; and (iv) whether a cadet’s wife originated from a Free State (with unmarried cadets coded as not having such a spouse). These data are obtained from Ancestry.com and FamilySearch.com, including linked data from the 1860 Census. We describe our data-linking procedure in the Supplemental Appendix A.1.
@@ 3.1 Descriptive Evidence
Note. This figure plots the relationship between the share of the slave population in a state and the likelihood that cadets from that state joined the Union army. Panel A concentrates on the states that participated in the war, while Panel B uses all cadets, including those who did not participate in the war. The circle’s size represents the number of West Point cadets hailing from each state.
@@ 3.2 Statistical Analysis
$X_{i,s}$ includes a range of personal characteristics. Our main analysis focuses on variables available for all cadets recorded in (Cullum, 1891): age in 1860, academic rank at West Point, and the slave population share in the cadet’s home state. When home-state fixed effects are included, state-level characteristics are naturally absorbed. In addition, we consider alternative measures of political-economic background, including whether a cadet owned slaves in 1860, and indicators for whether the cadet’s father, mother, or wife originated from Free States. These variables are constructed for the subset of cadets for whose information is available from Ancestry.com and FamilySearch.com.
@@
Our specification examines peer effects based on exogenous peer characteristics and not endogenous peer choices (Manski (1993)). Given that cadets’ decisions were made some years after graduation, it is less likely that there was direct coordination on decisions rather than lasting impacts of interactions from their formative years. Nonetheless, it is possible that there is correlation in their decisions, and thus for robustness, we also re-estimate the model using standard errors clustered at the cohort level as well as bootstrapped standard errors clustered at the cohort level (Supplemental Appendix Table S6). We also directly examine residual dependence across peer groups by estimating intra-class correlation coefficients (ICC in Supplemental Appendix Table S6) for the baseline residuals within cohorts. The results show negligible within-cohort residual correlation, providing no evidence of jointly determined wartime allegiance decisions.
@@ 3.3 The Impact of Peers on Cadets’ Allegiance
Note. This table presents the impact of the fraction of peers from Free States in a cadet’s cohort on that cadet’s decision to join the Union army. Columns (1)-(4) focus on cadets from Slave States (i.e., with a slave population higher than 1%) and Columns (5)-(8) on those from Free States (i.e., with a slave population lower than 1%). The variable Free-State father and Free-State mother equal 1 if a parent is from a Free State, and 0 otherwise. A subset of observations have missing information on parental origin. Free-State wife equals 1 if the wife is from a Free State and 0 otherwise. If the wife’s information is missing, we assign a value of 0. The standard errors presented in the parentheses are obtained through bootstrapping with 400 resampling iterations. ***$p<0.01$, **$p<0.05$, *$p<0.1$.
We include cadets who graduated between 1820 and 1860 in our baseline analysis because cadets dating back to the early cohorts (from 1820) participated in the Civil War, often serving in officer or administrative roles. As a robustness check, we re-estimate our baseline specifications after excluding those aged 60 (or 50) and above; the estimated peer effects remain quantitatively similar and statistically robust. The results are reported in Appendix Table S7.
@@
For Free-State cadets $S_{i}$ is 0 and so this probability is close to 1. For cadets heavily exposed to slavery (either from a region heavily dependent on slavery, or personally owning slaves), $S_{i}$ is large and this probability is close to 0. It is for intermediate values of $S_{i}-V$ that this varies substantially as a function of $\beta F_{i}$.
@@
We further divide the Slave States into groups by the relative slave populations. It turns out that the peer influence operates entirely on cadets from the range of Slave States that have slave populations below one third of the population. In Panel B of Table 2, we break states into three groups: those without slaves, those with slaves up to one third of the population, and those with slave populations above one third. We now see that the peer effect comes entirely from this middle group of states, and in fact the estimated coefficient and significance are larger than that from the previous table. These findings suggest that in High-Slavery States, economic imperatives and entrenched social norms overrode peer influence, muting its effect. This is consistent with a very high $S_{i}$ in equation (2).